A company remains an operating profit margin of 8% and sales-to-assets ratio (asset turnover ratio) of 3. It has assets of 2’000’000$ and equity of 1’200’000$. Its long term debt is 800’000$. Interest payments are 120’000$ and the tax rate is 35%.

  1. How much is sales?
  2. what is the ROA
  3. what is the ROE
  4. what is the ROC
Answer #1
Dear Student
Thank you for using homeworklib
Please find below the answer
Statementshowing Computations
Paticulars Amount
asset turnover ratio                            3.00
Assets           2,000,000.00
Sales = 2m*3           6,000,000.00
b)
Operating profit = 6m*8%               480,000.00
Less Interest               120,000.00
Income before tax               360,000.00
Tax at 35%            (126,000.00)
Income after tax               234,000.00
ROA = 234,000/2000,000 11.70%
c) ROE = 234,000/1200,000 19.50%
d) ROC = 234,000/(1200,000+800,000) 11.70%